# Positive operator-valued noncommutative polynomials are squares

### v1

@@ 4. Topology on $\mathcal{A}_{d}$
Let $(p_{\alpha})_{\alpha}$ be a net in $\mathcal{A}_{d}$ and $p\in\mathcal{A}_{d}.$ The sequence $(p_{\alpha})_{\alpha}$ converges to $p$ in the topology of $\mathcal{A}_{d}$ if and only if $(p_{\alpha}(A))$ converges to $p(A)$ in the weak operator topology. Here, $A$ is the tuple of operators defined in (3.2).
Lemmas 4.2 and 4.3, taken together, prove Proposition 4.1.
@@
If $(p_{\alpha}=\sum_{w}P_{\alpha,w}w)_{\alpha}\subset\mathcal{A}_{d}$ converges to $p=\sum_{w}P_{w}w\in\mathcal{A}_{d}$, then $(p_{\alpha}(X))_{\alpha}$ converges to $p(X)$ in the WOT for any $\tt{g}$}-tuple of self-adjoint operators $X=(X_{1},\dots,X_{{\tt{g}}}).$
@@
Fix a $\tt{g}$}}-tuple of self-adjoint operators $X=(X_{1},\dots,X_{{\tt{g}}}).$ By hypotheses, the nets $(P_{\alpha,w})_{\alpha}$ converge WOT for each $w.$ Thus, by Lemma A.1, for each fixed $w\in\langle x\rangle_{d},$ the net of operators $(P_{\alpha,w}\otimes X^{w})_{\alpha}$ converges to $P_{\alpha}\otimes X^{w}$ in the WOT, since $X^{w}$ is a fixed operator. Thus, $(p_{\alpha}(X)=\sum_{w}P_{\alpha,w}\otimes X^{w})_{\alpha}$ is a finite sum of convergent nets. ∎
Lemma 4.3.
Let $(p_{\alpha}=\sum_{w}P_{\alpha,w}w)_{\alpha}$ be a net in $\mathcal{A}_{d}$ and $A=(A_{1},\dots,A_{{\tt{g}}})$ be the tuple of self-adjoint operators of (3.2). If $(p_{\alpha}(A))_{\alpha}$ converges in the WOT, then $(p_{\alpha})_{\alpha}$ converges. Moreover, if $p=\sum_{w}P_{w}w\in\mathcal{A}_{d}$ is the limit of $(p_{\alpha}),$ then $p(A)$ is the limit of the convergent net $(p_{\alpha}(A))_{\alpha}.$
Proof.
Let $Z(p_{\alpha})$ be the column vector corresponding to $p_{\alpha}$ as in (3.3). Since $(p_{\alpha}(A))_{\alpha}$ converges in the WOT, $(Z(p_{\alpha}))_{\alpha}$ converges in the WOT. Now by (3.4), it follows that $(P_{\alpha,w})_{\alpha}$ converges in the WOT. Let $P_{w}$ be the WOT-limit of $(P_{\alpha,w})_{\alpha}.$ Then $p=\sum_{w}P_{w}w$ is the limit of $(p_{\alpha})_{\alpha}.$
Proposition 4.4.
The convex cone $\mathcal{C}_{d}$ is closed.
Proof.
Let $(p_{\alpha})_{\alpha}$ be a net in $\mathcal{C}_{d}$ that converges to $p\in\mathcal{A}_{2d}.$ Our aim is to show that $p\in\mathcal{C}_{d}.$ By Proposition 2.2, for each $\alpha$, there exists an $S_{\alpha}\in\mathcal{B}(\mathcal{H})^{N(d)\times N(d)}$ such that $S_{\alpha}\succeq 0$ and $V_{d}^{*}S_{\alpha}V_{d}=p_{\alpha}.$ Since the net of operators $(p_{\alpha}(A))_{\alpha}$ converges in the WOT, it follows from the Uniform Boundedness Principle that
@@
$\sup\limits_{\alpha}\|S_{\alpha}\|\ <\ \infty.$
@@
Thus, by the Banach–Alaoglu theorem, there is a sequence of operators $(S_{\alpha_{i}})_{i\in\mathbb{N}}\subset(S_{\alpha})_{\alpha}$ that converges to some operator $S\in\mathcal{B}(\mathcal{H})^{N(d)\times N(d)}$ in the ultraweak operator topology. The ultraweak operator topology is the weak-* topology on $\mathcal{B}(\mathcal{H})$ induced by the predual $\mathcal{T}(\mathcal{H})$, the trace class operators on $\mathcal{H}.$ It is the weakest topology such that predual elements remain continuous on $\mathcal{B}(\mathcal{H}).$ Since the convergence in the ultraweak operator topology is stronger than convergence in the weak operator topology, $S_{\alpha_{i}}\rightarrow S$ in the WOT. Thus, $S\succeq 0.$
Let $q=V_{d}^{*}SV_{d}\in\mathcal{C}_{d}$ and observe,
$p_{\alpha}(A)\ =\ V_{d}(A)^{*}(S_{\alpha}\otimes I_{\mathcal{F}^{2}_{{\tt{g}}}})V_{d}(A),$
where
$V_{d}(A):\ \mathcal{H}\otimes\mathcal{F}^{2}_{{\tt{g}}}\to\mathcal{H}^{N(d)}\otimes\mathcal{F}^{2}_{{\tt{g}}},\qquad h\otimes\xi\mapsto(h\otimes A^{w}\xi)_{w}.$
Since $(S_{\alpha_{i}})$ converges $S$ in the WOT,
$V_{d}(A)^{*}(S_{\alpha_{i}}\otimes I_{\mathcal{F}^{2}_{{\tt{g}}}})V_{d}(A)\ \to\ V_{d}(A)^{*}(S\otimes I_{\mathcal{F}^{2}_{{\tt{g}}}})V_{d}(A)=q(A)$
@@
in the WOT by Lemma A.1. Thus, $(p_{\alpha_{i}}(A))$ converges to $q(A)$ in the WOT. Therefore, $p(A)=q(A).$ Since, by the extraction formula, equation (3.4), the coefficients of $p$ and $q$ are determined by evaluation at $A,$ it follows that $p=q.$ Hence, $p\in\mathcal{C}_{d}.$
@@ 5. GNS Construction
In preparation of the application of the Hahn-Banach-convex separation theorem in the proof of Theorem 1.1 in Section 6, we establish a suitable version of the GNS construction in Proposition 5.3. The following result is well-known. We refer the reader to [Dav25, Section 3.1] and Exercise 3.4 in that book.
@@
Let $\mathcal{H}$ be a Hilbert space. If $f:\mathcal{B}(\mathcal{H})\to\mathbb{C}$ is a WOT continuous linear functional, then there exist a finite index set $J,$ vectors $h_{j},k_{j}\in\mathcal{H}$ such that
@@
If $\varphi:\mathcal{A}_{d}\to\mathbb{C}$ is a continuous linear functional, then there exist a finite index set $J,$ vectors $h_{j},k_{j}\in\mathcal{H}$ and scalars $c_{w}\in\mathbb{C}$ $(w\in\langle x\rangle_{d})$ such that for every $p=\sum P_{w}w\in\mathcal{A}_{d},$
@@
First identify $\mathcal{A}_{d}$ with $\mathcal{B}(\mathcal{H})^{N(d)}$ via
@@
This identification induces the product weak operator topology on $\mathcal{B}(\mathcal{H})^{N(d)},$ i.e., the topology generated by the seminorms $\|\cdot\|_{w,h,k}$ given by
$\|P\|_{w,h,k}\ :=\ |\langle P_{w}h,k\rangle|,\quad w\in\langle x\rangle_{d},\ h,k\in\mathcal{H},$
where $P=(P_{w})_{w}$ is a tuple of $N(d)$ many operators in $\mathcal{B}(\mathcal{H}).$ Now the statement follows from Lemma 5.1. ∎

### v2

@@ 4.1. Closedness of the cone $\mathcal{C}_{d}$
The convex cone $\mathcal{C}_{d}$ is closed in $\mathcal{A}_{\lcal{d}}$.
Firstly, since $\mathcal{A}_{2d}$ is closed in $\mathcal{A}_{\lcal{d}}$, it suffices to show $\mathcal{C}_{d}$ is closed in $\mathcal{A}_{2d}$ (in its product WOT topology). The proof of Proposition 4.1 proceeds in two steps. We first show that the cone $\mathcal{C}_{d}$ is closed with respect to the stronger ultraweak topology. Next, we prove that if a polynomial in $\mathcal{A}_{2d}$ can be separated from the cone $\mathcal{C}_{d}$ by an ultraweak continuous linear functional, then it can also be separated by a (possibly different) WOT continuous linear functional. This implication will yield the desired closedness of $\mathcal{C}_{d}$.
4.1.1. Closedness of $\mathcal{C}_{d}$ in the product ultraweak topology
Equip $\mathcal{A}_{\lcal{d}}$ with a norm in the following way. For $p=\sum\limits_{w\in\langle x\rangle_{\lcal{d}}}P_{w}w$,
$\|p\|\ =\ \sum\limits_{w\in\langle x\rangle_{\lcal{d}}}\|P_{w}\|,$
where $\|T\|$ denotes the operator norm for $T\in\mathcal{B}(\mathcal{H}).$
@@ 4.1.1. Closedness of $\mathcal{C}_{d}$ in the product ultraweak topology
For any $t>0$, the truncated cone
$\mathcal{C}_{d,t}\ :=\ \{p\in\mathcal{C}_{d}\ :\ \|p\|\leq t\}$
is closed in the ultraweak topology on $\mathcal{A}_{\lcal{d}}.$
@@
Fix $t>0.$ Let $(p_{\alpha})_{\alpha}$ be a net in $\mathcal{C}_{d,t}$ that converges to $p\in\mathcal{A}_{2d}.$ Our aim is to show that $p\in\mathcal{C}_{d,t}.$ Note that for any $\alpha$,
$\|p_{\alpha}(A)\|\ \leq\ \|p_{\alpha}\|\,\|A\|\ \leq\ t\|A\|.$
Thus,
@@
$\sup\limits_{\alpha}\|G_{\alpha}\|\ <\ \infty$
@@
for any $G_{\alpha}\in\Gamma_{p_{\alpha}}$. Thus, by the Banach–Alaoglu theorem, there is a subnet of operators $(G_{\beta})_{\beta}$ that converges to some operator $G\in\mathcal{B}(\mathcal{H})^{N(d)\times N(d)}$ in the ultraweak topology. Since the convergence in the ultraweak operator topology is stronger than convergence in the weak operator topology, $G_{\beta}\longrightarrow G$ in the WOT. Thus, $G\succeq 0.$
Let $q=V_{d}^{*}GV_{d}\in\mathcal{C}_{d}$ and observe,
$p_{\alpha}(A)\ =\ V_{d}(A)^{*}(G_{\alpha}\otimes I_{\mathcal{F}^{2}_{{\tt{g}}}})V_{d}(A),$
where $V_{d}(A):\ \mathcal{H}\otimes\mathcal{F}^{2}_{{\tt{g}},\lcal{d}}\to\mathcal{H}^{N(d)}\otimes\mathcal{F}^{2}_{{\tt{g}},\lcal{d}},$ is defined by
$V_{d}(A)h\otimes\xi=(h\otimes A^{w}\xi)_{w}.$
Since $(G_{\beta})_{\beta}$ converges to $G$ in the WOT and the tuple $A$ acts on a finite dimensional Hilbert space,
$V_{d}(A)^{*}(G_{\beta}\otimes I_{\mathcal{F}^{2}_{{\tt{g}}}})V_{d}(A)\ \longrightarrow\ V_{d}(A)^{*}(G\otimes I_{\mathcal{F}^{2}_{{\tt{g}}}})V_{d}(A)\ =\ q(A)$
@@
in the WOT. Thus, $(p_{\beta}(A))_{\beta}$ converges to $q(A)$ in the WOT. Therefore, $p(A)=q(A).$ Since, by the extraction formula, equation (3.6), the coefficients of $p$ and $q$ are determined by evaluation at $A,$ it follows that $p=q.$ Hence, $p\in\mathcal{C}_{d}.$ Since the norm is lower semicontinuous in the ultraweak topology, $\|p\|\leq t.$ Hence, $p\in\mathcal{C}_{d,t}.$
It follows from Krein–Smulian Theorem (see, e.g., [Dav25, Theorem 3.6.2]) that the cone $\mathcal{C}_{d}$ is closed in the ultraweak topology.
4.1.2. Closedness of $\mathcal{C}_{d}$ in the product WOT topology
Lemma 4.3.
Given trace-class operators $S_{i}\in\mathcal{T}(\mathcal{H})$, $1\leq i\leq m$ and $\epsilon>0$, there exists a finite-rank operator $P$ such that
$\|S_{i}-PS_{i}P\|_{1}<\epsilon$
for all $i$. Here $\|\cdot\|_{1}$ denotes the trace norm.
Proof.
Since finite-rank operators are dense in $\mathcal{T}(\mathcal{H})$ with respect to the trace norm, there are finite-rank operators $F_{i}$ with
$\|S_{i}-F_{i}\|_{1}<\frac{\epsilon}{2}.$
Let $P$ be the orthogonal projection onto $\sum_{i}{\rm ran}(F_{i})+\sum_{i}{\rm ran}(F_{i}^{*})$. Then $PF_{i}P=F_{i}$ for all $i$. Hence
$\begin{split}\|S_{i}-PS_{i}P\|_{1}&\leq\|S_{i}-F_{i}\|_{1}+\|F_{i}-PS_{i}P\|_{1}\\ &=\|S_{i}-F_{i}\|_{1}+\|PF_{i}P-PS_{i}P\|_{1}\\ &=\|S_{i}-F_{i}\|_{1}+\|P(F_{i}-S_{i})P\|_{1}\\ &\leq\frac{\epsilon}{2}+\|P\|\,\|F_{i}-S_{i}\|_{1}\,\|P\|\\ &\leq\frac{\epsilon}{2}+\frac{\epsilon}{2}=\epsilon.\qed\end{split}$
Lemma 4.4.
If $\varphi:\mathcal{A}_{2d}\to\mathbb{C}$ be an ultraweak continuous linear functional that separates the cone $\mathcal{C}_{d}$ from a fixed polynomial $p$ in $\mathcal{A}_{2d},$ that is,
$\varphi(r^{*}r)\geq 0\quad\text{for all $r\in\mathcal{A}_{d}$}\quad\text{and}\quad\varphi(p+p^{*})<0,$
then there exists a WOT continuous linear functionl $\tilde{\varphi}:\mathcal{A}_{2d}\to\mathbb{C}$ such that
$\tilde{\varphi}(r^{*}r)\geq 0\quad\text{for all $r\in\mathcal{A}_{d}$}\quad\text{and}\quad\tilde{\varphi}(p+p^{*})<0.$
Proof.
Since $\varphi$ is ultraweak continuous, there exist trace class operators $S_{w}$ $(w\in\langle x\rangle_{2d})$ in $\mathcal{B}(\mathcal{H})$ such that
$\varphi(q)=\sum\limits_{w\in\langle x\rangle_{2d}}{\rm Tr}\,(S_{w}Q_{w}),$
where $q=\sum_{w\in\langle x\rangle_{2d}}Q_{w}w.$ For $r,r^{\prime}\in\mathcal{A}_{d}$,
$\varphi(r^{*}r^{\prime})\ =\ \sum\limits_{u,v\in\langle x\rangle_{d}}{\rm Tr}\,(S_{u^{*}v}R_{u}^{*}R^{\prime}_{v}),$
where $r=\sum_{u\in\langle x\rangle_{d}}R_{u}u$ and $r^{\prime}=\sum_{v\in\langle x\rangle_{d}}R^{\prime}_{v}v.$ Denote by $S$ the $N(d)\times N(d)$ block operator matrix whose $(u,v)$ entry is $S_{v^{*}u}.$
For any $r\in\mathcal{A}_{d}$, define the row operator $R:\bigoplus_{u\in\langle x\rangle_{d}}\mathcal{H}\to\mathcal{H}$ by
$R(\oplus_{u\in\langle x\rangle_{d}}\zeta_{w})\ =\ \sum_{u}R_{u}\zeta_{u}.$
Thus,
$\varphi(r^{*}r)\$$=\ \sum\limits_{u,v\in\langle x\rangle_{d}}{\rm Tr}\,(S_{u^{*}v}R_{u}^{*}R_{v})$
$=\ \sum_{u,v\in\langle x\rangle_{d}}{\rm Tr}\,\left([S]_{v,u}[R^{*}R]_{u,v}\right)$
$=\ \ \ \ {\rm Tr}\,(SR^{*}R).$
We claim that
${\rm Tr}\,(ST)\ \geq\ 0$
for any positive operator $T\in\mathcal{B}(\bigoplus_{\langle x\rangle_{d}}\mathcal{H}).$ It follows from Lemma 2.1 that $T=R^{*}R$ for some $R:\bigoplus_{\langle x\rangle_{d}}\mathcal{H}\to\mathcal{H}.$ Letting $r=\sum R_{u}u,$ where $R_{u}$ is the $u^{\rm th}$ element of the row operator $R$ gives
${\rm Tr}\,(ST)\ =\ {\rm Tr}\,(SR^{*}R)\ =\ \varphi(r^{*}r)\ \geq\ 0.$
To prove that $S$ is a positive operator, it remains to show that $S$ is self-adjoint. For $T\succeq 0,$ ${\rm Tr}\,(ST)$ is real. Hence,
${\rm Tr}\,(S^{*}T)\ =\ {\rm Tr}\,(TS^{*})\ =\ {\rm Tr}\,((ST)^{*})\ =\ \overline{{\rm Tr}\,(ST)}\ =\ {\rm Tr}\,(ST).$
${\rm Tr}\,(S^{*}T)\ =\ {\rm Tr}\,(ST)\quad\text{for all $T\in\mathcal{B}(\bigoplus_{\langle x\rangle_{d}}\mathcal{H})$ }.$
Hence $S^{*}=S$ and $S\succeq 0.$
The finitely many trace class operators $S_{w}$, $w\in\langle x\rangle_{2d},$ can be approximated by finite rank operators in the trace norm as in Lemma 4.3. That is, for any $n\in\mathbb{N}$, there exists a finite-rank projection $P_{n}$ of $\mathcal{H}$ such that
$\|S_{w}-P_{n}S_{w}P_{n}\|_{1}<\frac{1}{n},$
for all $w\in\langle x\rangle_{2d},$ where $\|\cdot\|_{1}$ denotes the trace norm. Letting $S^{(n)}$ denote the $N(d)\times N(d)$ block operator matrix whose $(u,v)$ entry is $P_{n}S_{v^{*}u}P_{n},$
$S^{(n)}\ =\ (I_{{N(d)}}\otimes P_{n})\,S\,(I_{{N(d)}}\otimes P_{n}).$
Whence $S^{(n)}$ is a finite-rank psd operator.
Define a linear functional $\varphi_{n}:\mathcal{A}_{2d}\to\mathbb{C}$ by
$\varphi_{n}(Qu^{*}v)={\rm Tr}\,(P_{n}S_{u^{*}v}P_{n}\ Q)$
for $Q\in\mathcal{B}(\mathcal{H})$ and $u,v\in\langle x\rangle_{d}$. For $r=\sum_{u\in\langle x\rangle_{d}}R_{u}u$ and $r^{\prime}=\sum_{v\in\langle x\rangle_{d}}R^{\prime}_{v}v,$
$\begin{split}\varphi_{n}(r^{*}r^{\prime})&=\sum_{u,v\in\langle x\rangle_{d}}\varphi_{n}(R_{u}^{*}R^{\prime}_{v}u^{*}v)=\sum_{u,v\in\langle x\rangle_{d}}{\rm Tr}\,(P_{n}S_{u^{*}v}P_{n}\ R_{u}^{*}R^{\prime}_{v})\\ &={\rm Tr}\,(P_{n}SP_{n}\ R^{*}R^{\prime})={\rm Tr}\,(S^{(n)}\ R^{*}R^{\prime}).\end{split}$
Since $S^{(n)}$ is psd,
$\varphi_{n}(r^{*}r)\ =\ {\rm Tr}\,(S^{(n)}R^{*}R)\geq 0$
for all $r\in\mathcal{A}_{d}.$ Further, since $S^{(n)}$ is a finite-rank operator, $\varphi_{n}$ is WOT continuous.
Since $(P_{n}S_{u^{*}v}P_{n})_{n}$ converges to $S_{u^{*}v}$ in the trace norm, ${\rm Tr}(P_{n}S_{u^{*}v}P_{n}Q)$ converges to ${\rm Tr}(SQ).$ Hence $\varphi_{n}(p+p^{*})$ converges to $\varphi(p+p^{*}).$ As $\varphi(p+p^{*})<0$, there exists a natural number $n$ such that $\varphi_{n}(p+p^{*})<0.$ The linear functional $\tilde{\varphi}:=\varphi_{n}$ has the desired separation properties. ∎
We are finally ready to prove Proposition 4.1.
Proof of Proposition 4.1.
Suppose $p\in\mathcal{A}_{2d}$ is not in $\mathcal{C}_{d}$. Since $\mathcal{C}_{d}$ is ultraweak closed, there is an ultraweak continuous linear functional $\varphi$ on $\mathcal{A}_{2d}$ that separates $p$ from $\mathcal{C}_{d}$. Now apply Lemma 4.4 to obtain a WOT continuous separating linear functional $\tilde{\varphi}$. Thus $p$ is not in the WOT closure of $\mathcal{C}_{d}$. ∎
@@ 5. GNS Construction
In preparation of the application of the Hahn-Banach-convex separation theorem in the proof of Theorem 1.1 in Section 6, we establish a suitable version of the GNS construction in Proposition 5.4. Before doing so, for the reader’s convenience, we state three well-known lemmas that will be used in the proof.
@@
Let $\mathcal{H}$ be a Hilbert space. If $f:\mathcal{B}(\mathcal{H})\to\mathbb{C}$ is a WOT continuous linear functional, then there exist a finite index set $J$ and vectors $h_{j},k_{j}\in\mathcal{H}$ such that
@@
If $\varphi:\mathcal{A}_{\lcal{d}}\to\mathbb{C}$ is a continuous linear functional, then there exist a finite index set $J,$ vectors $h_{j},k_{j}\in\mathcal{H}$ and scalars $c_{w}\in\mathbb{C}$ $(w\in\langle x\rangle_{\lcal{d}})$ such that for every $p=\sum P_{w}w\in\mathcal{A}_{\lcal{d}},$
@@
First identify $\mathcal{A}_{\lcal{d}}$ with $\mathcal{B}(\mathcal{H})^{N(\lcal{d})}$ via
@@
This identification induces the product weak operator topology on $\mathcal{B}(\mathcal{H})^{N(\lcal{d})},$ i.e., the topology generated by the seminorms $\|\cdot\|_{w,h,k}$ given by
$\|P\|_{w,h,k}\ :=\ |\langle P_{w}h,k\rangle|,\quad w\in\langle x\rangle_{\lcal{d}},\ h,k\in\mathcal{H},$
where $P=(P_{w})_{w}$ is a tuple of $N(\lcal{d})$ many operators in $\mathcal{B}(\mathcal{H}).$ Now the statement follows from Lemma 5.1. ∎
The third lemma presents a few basic properties of the vectorization map used in the proof of Proposition 5.4. See Subsection 5.5. Let $HS(\mathcal{H},\mathcal{E})$ denote the Hilbert-Schmidt operators from the Hilbert space $\mathcal{H}$ to the Hilbert space $\mathcal{E}.$ For a fixed orthonormal basis $(e_{\delta})_{\delta}$ of $\mathcal{H}$, the vectorization map, $\mathrm{vec}:HS(\mathcal{H},\mathcal{E})\to\mathcal{H}\otimes\mathcal{E}$ is defined, for $T\in HS(\mathcal{H},\mathcal{E}),$ by
$\mathrm{vec}(T)\ =\ \sum_{\delta}e_{\delta}\otimes Te_{\delta}$
Lemma 5.3.
Let $A,B\in{HS}(\mathcal{H},\mathcal{E})$ and $P,Q\in\mathcal{B}(\mathcal{H})$, $T\in\mathcal{B}(\mathcal{E})$.
  1. (1)
    Independent of the choice of orthonormal basis,
    In particular, $\|\mathrm{vec}(A)\|=\|A\|_{\mathrm{HS}}.$
  2. (2)
    The following compatibility with left/right actions holds:
    hence, more generally,
    $(P\otimes T)\,\mathrm{vec}(A)=\mathrm{vec}(T\,A\,P^{*}).$
  3. (3)
    (Inner–product identity used in Proposition 5.4)
Proof.
Fix an orthonormal basis $(e_{\delta})_{\delta}$ of $\mathcal{H}$.
Using the definition of $\mathrm{vec},$
The right-hand side is independent of the choice of orthonormal basis, hence so is the left-hand side, proving item 1.
To prove the first identity of item 2, observe
Turning to item 3, using item 2 twice and then item 1,
We have now reached the main result of this section.